OpenStudy (anonymous):

A certain weak acid, HA, has a Ka value of 1.8×10−7. Part A Calculate the percent dissociation of HA in a 0.10 M solution.

OpenStudy (anonymous):

Give me a second to do this..

OpenStudy (anonymous):

Dissociation equation of the acid; HA(aq) <-----> H+(aq) + A-(aq) 0.010-x M ........ x M ........ x M Ka = [H+] [A-] / [HA] 1.3x10^-7 = (x)(x) / (0.010 - x) Since the amount of x is very small compared to 0.010 M, it is neglected. 1.3x10^-7 = (x)(x) / (0.010) x^2 = 1.3x10^-9 x = 3.6x10^-5 M Since 3.6x10^-5 M is the amount dissociared (ionized) and 0.010 M is the initial amount, the percent ionization of the acid will be: (3.6x10^-5 M / 0.010 M) x 100 = 0.36 %

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

thanks !